# The Developing Human Connectome Project: A Fast Deep Learning-based Pipeline for Neonatal Cortical Surface Reconstruction

### v1

@@ 4.1.3 Surface Topology
We further examine surface self-intersections, which usually appear in the narrow sulci of pial surfaces and will affect the estimation of cortical morphological features [38]. We detect and count the number of self-intersecting faces (SIFs) in white matter and pial surface meshes generated by both original and DL-based dHCP pipelines. To verify that the proposed multiscale deformation networks can effectively reduce SIFs, we conduct ablation experiments by removing the Gaussian smoothing layer or replacing the scaling and squaring method [2] with the forward Euler method for integration. The number of SIFs are reported in Table 4. It shows that our DL-based dHCP pipeline produces only 0.18 SIFs on average in white matter surfaces. The original dHCP pipeline produces marginally fewer SIFs in pial surfaces than ours since it performs collision detection during the iterative surface deformation. Table 4 also validates that SIFs can be reduced effectively by adding the Gaussian smoothing layer and using the scaling and squaring method.
4.1.4 Generalization ability
In addition to the dHCP neonatal dataset [17], we evaluate the generalization ability of our DL-based dHCP pipeline on an unseen dataset. We consider the Evaluation of Preterm Imaging (ePrime) dataset [^3] https://www.npeu.ox.ac.uk/research/projects/22-eprime-mr-imaging. [16, 27], comprising 486 T2w MRI scans of neonates with gestational age between 23 and 33 weeks, in which 444 subjects are scanned at term-equivalent age between 38 and 45 weeks. The MR images are acquired with a Philips Intera 3T scanner, using a T2w turbo echo sequence (TES) with parameters repetition/echo time TR/TE=8670/160ms and TSE factor of 16. The ePrime dataset has relatively low image resolution with the voxel size of $0.86\times 0.86\times 1.0mm^{3}$, and thus it is challenging to extract high-quality cortical surfaces.
To test the capacity of generalization, we run both original and our DL-based dHCP pipelines on unseen T2w MRI scans of the ePrime dataset without fine-tuning. The qualitative comparison in Figure 12 shows that there are severe corruptions and discontinuity on the gyri of cortical surfaces generated by the original dHCP pipeline, while our DL-based pipeline produces cortical surfaces with acceptable quality. This demonstrates that our DL-based pipeline has better generalize ability on the unseen and low-resolution dataset.
@@ 4.2 Sphere Quality
Compared to [70], our approach has several advantages. First, [70] needs to inflate and project the white matter surfaces to create initial spheres, which requires 1.79 seconds for each brain hemisphere with GPU acceleration in our experiments. Instead, our approach uses an fixed initial sphere for all subjects without extra processing. The initial sphere is generated by FreeSurfer with minimized metric distortions as reported in Table 5. This provides a strong prior with small initial distortions for the training. Such a global initial surface also saves the runtime of interpolation and resampling, since the barycentric coordinates can be pre-computed. Hence, our approach is capable of projecting the cortical surface to a sphere in 0.33 seconds, which is 6$\times$ faster than [70]. Furthermore, [70] computes the metric distortions between the resampled white matter surface and the icosphere, and thus requires multiscale networks to improve the performance. As reported in Table 5, [70] does not outperform FreeSurfer in the case of single-scale architecture. Our end-to-end metric distortion loss is more effective so that a single-scale Spherical U-Net is sufficient to produce significantly smaller distortions than FreeSurfer.
@@ 5 Discussion
The quality control results validate that for more than 80% of test samples, our DL-based dHCP pipeline produces more anatomically accurate (54.2%) or equal quality (28.3%) cortical surfaces compared to the original dHCP pipeline [44]. Particularly, the visual inspection shows that our learning-based cortical surface reconstruction approach can effectively reduces the artifacts on the top and occipital lobe of the cortical surfaces produced by original dHCP pipeline. We also verify that the proposed DL-based dHCP pipeline is capable of generalizing well on unseen and low-resolution ePrime neonatal brain MRI data [16, 27]. In addition, our unsupervised learning-based spherical projection framework significantly reduces both edge and area distortions compared to previous work while only requires 0.33 seconds of runtime for each brain hemisphere.
However, despite the high efficacy and efficiency of the proposed dHCP DL pipeline, the anatomical accuracy of predicted cortical surfaces is inevitably restricted by the quality of pseudo GT cortical surfaces generated by the original dHCP pipeline, due to the limitations of supervised learning. Besides, the proposed multiscale deformation network predicts cortical surfaces from the input T2w brain MRI directly without the need of segmentations. Such an end-to-end framework facilitates the surface reconstruction procedure and avoids the potential errors in cortical surfaces introduced by imprecise segmentations, while this also means that our dHCP DL pipeline does not provide brain tissue segmentations and cortical surface parcellation.
In future work, we will consider semi-supervised learning technique to further enhance the anatomical accuracy of the cortical surfaces. In addition to pseudo GT cortical surfaces, the unsupervised brain MRI intensity will be utilized to deform the cortical surfaces towards the WM/cGM or cGM/CSF boundary where the MRI achieves maximum image contrast. Furthermore, we plan to enrich the functionality and incorporate more GPU-accelerated processing steps into our DL-based dHCP pipeline, e.g., brain regional and tissue segmentations, cortical parcellation, and local gyrification index estimation.

### v2

@@ 4.1.4 Surface Topology
We further examine surface self-intersections, which usually appear in the narrow sulci of pial surfaces and will affect the estimation of cortical morphological features [44]. We detect and count the number of self-intersecting faces (SIFs) in white matter and pial surface meshes generated by both original and DL-based dHCP pipelines. To verify that the proposed multiscale deformation networks can effectively reduce SIFs, we conduct ablation experiments by removing the Gaussian smoothing layer or replacing the scaling and squaring method [33, 2] with the forward Euler for integration. The number of SIFs are reported in Table 4.
Although our multiscale deformation networks theoretically guarantee the diffeomorphisms, there are still negligible number of SIFs caused by numerical errors in the discretization, e.g., the ODE integration, the volumetric representation of deformation fields, and the limited number of mesh vertices. Table 4 shows that our DL-based dHCP pipeline produces only 0.18 SIFs on average in white matter surfaces. The original dHCP pipeline produces marginally fewer SIFs in pial surfaces than ours since it performs collision detection during the iterative surface deformation. Table 4 also validates that SIFs can be reduced effectively by adding the Gaussian smoothing layer and using the scaling and squaring method.
4.1.5 Generalizability
In addition to the dHCP neonatal dataset [18], we evaluate the generalizability of our DL-based dHCP pipeline on an unseen dataset. We consider the Evaluation of Preterm Imaging (ePrime) dataset [17, 30], comprising 486 T2w MRI scans of neonates with gestational age between 23 and 33 weeks, in which 444 subjects are scanned at term-equivalent age between 38 and 45 weeks. The MR images are acquired with a Philips Intera 3T scanner, using a T2w turbo echo sequence (TES) with parameters repetition/echo time TR/TE=8670/160ms and TSE factor of 16. The ePrime dataset has relatively low image resolution with the voxel size of $0.86\times 0.86\times 1.0mm^{3}$, and thus it is challenging to extract high-quality cortical surfaces.
To test the generalizability, we run both original and our DL-based dHCP pipelines on unseen T2w MRI scans of the ePrime dataset without fine-tuning. The qualitative comparison in Figure 15 shows that there are severe corruptions and discontinuity on the gyri of cortical surfaces generated by the original dHCP pipeline, while our DL-based pipeline produces cortical surfaces with acceptable quality. This demonstrates that our DL-based pipeline has better generalizability on the unseen and low-resolution dataset.
@@ 4.2 Sphere Quality
Compared to [77], our approach has several advantages. First, [77] needs to inflate and project the white matter surfaces to create initial spheres, which requires 1.79 seconds for each brain hemisphere with GPU acceleration in our experiments. Instead, our approach uses a fixed initial sphere for all subjects without extra processing. The initial sphere is generated by FreeSurfer with minimized metric distortions as reported in Table 5. This provides a strong prior with small initial distortions for the training. Such a global initial surface also saves the runtime of interpolation and resampling, since the barycentric coordinates can be pre-computed. Hence, our approach is capable of projecting the cortical surface to a sphere in 0.33 seconds, which is 6$\times$ faster than [77]. Furthermore, [77] computes the metric distortions between the resampled white matter surface and the icosphere, and thus requires multiscale networks to improve the performance. As reported in Table 5, [77] does not outperform FreeSurfer in the case of single-scale architecture. Our end-to-end metric distortion loss is more effective so that a single-scale Spherical U-Net is sufficient to produce significantly smaller distortions than FreeSurfer.
@@ 5 Discussion
The results of visual surface quality control validate that for more than 80% of test samples, our DL-based dHCP pipeline produces cortical surfaces with higher (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline [50]. Particularly, the visual inspection on cortical surfaces and quantitative evaluation on surface distance demonstrate that our learning-based cortical surface reconstruction approach can effectively reduce the artifacts on the top and occipital lobe of the cortical surfaces produced by the original dHCP pipeline. We also verify that the proposed DL-based dHCP pipeline is capable of generalizing well on unseen and low-resolution ePrime neonatal brain MRI data [17, 30]. In addition, our unsupervised learning-based spherical projection framework significantly reduces both edge and area distortions compared to previous work while only requires 0.33 seconds of runtime for each brain hemisphere.
However, despite the high efficacy and efficiency of the proposed dHCP DL pipeline, the anatomical correctness of predicted cortical surfaces is inevitably restricted by the surface quality of pseudo GT cortical surfaces generated by original dHCP pipeline, due to the limitations of supervised learning. Besides, the proposed multiscale deformation network predicts cortical surfaces from the input T2w brain MRI directly without the reliance on segmentations. Such an end-to-end framework facilitates the surface reconstruction procedure and avoids the potential errors in cortical surfaces introduced by imprecise segmentations, while this also means that our DL-based dHCP pipeline does not provide brain tissue segmentations and cortical surface parcellation.
In future work, we will consider semi-supervised learning technique to further enhance the anatomical correctness of the cortical surfaces. In addition to pseudo GT cortical surfaces, the unsupervised brain MRI intensity will be utilized to deform the cortical surfaces towards the WM/cGM or cGM/CSF boundary where the MRI achieves maximum image contrast. Furthermore, we plan to enrich the functionality and incorporate more GPU-accelerated processing steps into our DL-based dHCP pipeline, e.g., brain regional and tissue segmentations, cortical parcellation, and local gyrification index estimation.